# KCL-PH-TH/2025-33

### v1

@@ IV Implications for Cosmology
$N_{\rm kin}\gtrsim 16.46+\log\left(\frac{10^{14}\mathrm{GeV}}{H_{\rm end}}\right)\,,$ (28)
@@
$\Delta^{2}_{\mathcal{R}}(k)=\Delta^{2}_{\mathcal{R},\mathrm{CMB}}(k)+\Delta^{2}_{\mathcal{R},\mathrm{peak}}\delta(k-k_{\rm PBH})\,.$ (29)
@@
$\beta\sim\frac{\Delta_{\mathcal{R},\mathrm{peak}}}{\sqrt{2\pi}\delta_{c}}\exp(-\frac{\delta_{c}^{2}}{2\Delta^{2}_{\mathcal{R},\mathrm{peak}}})\,.$ (30)
@@
$\beta\gtrsim\left(\frac{\Gamma_{\rm ev}}{H_{\rm end}}\right)^{1/2}e^{-\frac{3}{2}N_{\rm kin}+3(N_{\rm form}-N_{\rm end})}\,,$ (31)
@@
In Figure 6, we translate the results obtained in Section III for $\delta_{c}$ into lower bounds on $\Delta^{2}_{\mathcal{R},\mathrm{peak}}$ that satisfy the condition of Eq. (31) for an arbitrary value of $H_{\rm end}=10^{13}\mathrm{GeV}$. Nevertheless, due to the logarithmic dependency of $\Delta^{2}_{\mathcal{R},\mathrm{peak}}$ on $\beta$, we note that the results depend only very mildly on the actual value of $H_{\rm end}$. From the figure, it is clear that the very small values of $\delta_{c}$ we obtained for large perturbations (which form at larger $N_{\rm form}$) reduce the required amplitude of the power spectrum by more than an order of magnitude compared to previous claims in the literature that use the analytical results of [27].
This result is particularly important, as it is generally very challenging–from a model-building perspective–to enhance the power spectrum of comoving scalar curvature perturbations all the way up to $)\Delta^{2}_{\mathcal{R},\mathrm{peak}}\sim\order{10^{-2}\sim 10^{-3}}$ without requiring significant fine-tuning of the inflaton potential parameters [43]. Moreover, such enhancements often risk conflicting with the CMB predictions, especially regarding the spectral index $n_{s}$ and the tensor-to-scalar ratio $r$. However, according to our results, the required power spectrum amplitude during a kination era for PBHs to play a significant cosmological role only needs to reach ($)\sim\order{10^{-5}}$. Achieving this level of enhancement within an inflationary scenario featuring a transient ultra-slow-roll phase does not suffer from the same fine-tuning issues, and is therefore more plausible from an ultraviolet complete theoretical framework.
@@
V Summary and outlook
We have studied the evolution of large scalar inhomogeneities during a kination epoch beyond the linear regime using numerical relativity.
@@ V Summary and outlook
Starting with small perturbations, our results in the sub-horizon limit agree with perturbative analysis: linear perturbations grow relative to the background and eventually dominate the universe, but do not collapse into black holes. We then examined initially large perturbations and found that they rapidly begin to behave like radiation unless the initial non-linear parameter is big $b\gg 1$, in which case gravitational forces are strong enough to induce immediate collapse into a black hole. However, such large sub-horizon perturbations cannot arise from the growth of initially small fluctuations after horizon re-entry: their domination over the background would halt further growth, preventing collapse. We therefore conclude that only mechanisms capable of sourcing large sub-horizon inhomogeneities at early times could lead to gravitational collapse.
@@
On the other hand, super-horizon inhomogeneities manifest a much more interesting phenomenology with regard to strong gravity effects. The very long oscillation timescales of these modes imply that they behave like curvature perturbations prior to horizon re-entry. We observed that even in the limit where $b\ll 1$, such perturbations grow significantly relative to the background within just a few e-folds of kination. After horizon re-entry, their fate depends on their amplitude, they may either evolve into a radiation-like background–similar to typical for sub-horizon modes–or collapse into black holes. Our key finding is that the critical overdensity extrapolated from our numerical evolution, is much smaller compared to the perturbative predictions, although it converges to the perturbative predictions when the initial wavelength approaches $H_{0}^{-1}$. For increasingly larger wavelengths, we find that the extrapolated critical overdensity decreases, reaching around $\delta_{c,\text{numerical}}\approx 0.02$ at $\lambda_{0}\approx 2000H_{0}^{-1}$, as shown in Figure 4.
@@
This significant reduction of the extrapolated critical overdensity for the collapse of super-horizon modes into PBHs as compared to previously existing results in the literature is crucial. Indeed, the capacity of Gaussian perturbations—characterized by a primordial power spectrum $\Delta_{\mathcal{R}}^{2}$–to produce PBHs from their horizon re-entry is exponentially sensitive to this critical value, $\beta\simeq\exp(-\delta_{c}^{2}/\Delta_{\mathcal{R}}^{2})$. Reducing $\delta_{c}$ by over an order of magnitude thus considerably eases PBH formation. We estimated the typical value that would be required for $\Delta_{\mathcal{R}}^{2}$ for PBH formation during kination to lead to PBH domination and evaporation before BBN, and found that values of ($)\order{10^{-5}}$ were sufficient–compared to around ($)\order{10^{-3}}$ in the standard kination scenario.
@@
It is also worth noting that in our numerical simulations, the requirement that the initial spatial hyperslice be conformally flat, combined with the imposition of periodic boundary conditions, enforces a high degree of symmetry in our physical setup. Intuitively, one might expect that in the case of black hole formation, the collapse of a single perturbation mode in three spatial directions might remove some energy from the system to create the spherical configuration that eventually collapses. Therefore, we cannot exclude the scenario in which the numerical value of $\delta_{c}$ in the case of direct spherical collapse might be even lower than what we have inferred in section III. It is important that future simulations explore such spherically symmetric configurations so that the expression for $\delta_{c}$ can be verified more accurately; in this work, we did not pursue this direction due to the difficulty in finding initial conditions that both respect spherical symmetry and satisfy the integrability conditions imposed by periodic boundary conditions. On the opposite, additional degrees of non-sphericity might as well render the collapse more difficult to achieve [44], which also requires a specific study in the context of kination.
@@
Finally, it is necessary to comment on the choice of gauge. In our numerical simulations, we employ a modified version of the moving puncture gauge, which includes a contribution from the volume-averaged extrinsic curvature scalar. As explained in Section III, this particular gauge choice leads to the lapse function eventually dropping to very low values, effectively freezing the simulations for very large initial scalar wavelengths. A more suitable gauge could enable the evolution of the system for at least seven e-folds of kination, potentially allowing for a more accurate determination of $\delta_{c}$ in the limit of very large initial wavelengths. We leave such explorations to future work.
VI Acknowledgements
We acknowledge useful conversations with Martin Mosny, Joseph Conlon, Benjamin Stefanek, and Katy Clough. CC and PG thank Santiago Agüí Salcedo, Gonzalo Villa, Lucas Vicente Garcia-Consuegra, Josu Calvo Aurrekoetxea, Daniel G. Figueroa and Liina Chung-Jukko for their useful input. We would also like to thank the GRChombo team (http://www.grchombo.org/). This work is supported by a Research Project Grant RPG-2021-423 from Leverhulme Trust.
@@ Appendix A Linear Perturbations During Kination
We briefly review the relevant solutions of kination in cosmological perturbation theory context; for a more detailed analysis, we refer the reader to [17, 23, 24]. The scalar field in the presence of small scalar perturbations is expressed as $\phi\rightarrow\bar{\phi}(t)+\delta\phi(t,\textbf{x})$ with the perturbation decomposed into Fourier modes as
$\delta\phi(t,\textbf{x})=\int\frac{d^{3}\textbf{k}}{(2\pi)^{3}}\delta\phi_{\textbf{k}}e^{i\textbf{k}\cdot\textbf{x}}\,.$ (32)
In Newtonian gauge the metric reads
[…]
@@
In this case the equation of motion for the background scalar field $\bar{\phi}$ and its perturbation in Fourier space is given by:
$\ddot{\bar{\phi}}+3H\dot{\bar{\phi}}=0\,,$ (34)
$\ddot{\delta\phi}_{\textbf{k}}+3H\dot{\delta\phi}_{\textbf{k}}+\frac{\textbf{k}^{2}}{a^{2}}\delta\phi_{\textbf{k}}=4\dot{\Phi}_{\textbf{k}}\dot{\bar{\phi}}\,,$ (35)
with the source term on the right-hand side and the perturbation equation of motion given by the coupling of the Newtonian potential $\Phi_{\textbf{k}}$ and the $\delta\phi_{\textbf{k}}$ perturbation. The scale factor evolves as
$a=a_{0}(t/t_{0})^{c}$
with $c=2/3(1+w)\equiv 1/3$ for kination and the Hubble parameter $H=\sqrt{8\pi G/3\bar{\rho}}=c/t$. To fully obtain the solution in the perturbation theory, the system of those PDEs is closed by the Poisson’s equation for the Newtonian potential $\Phi$
$\nabla^{2}\Phi_{\textbf{k}}=-4\pi Ga^{2}(\rho_{\textbf{k}}-\bar{\rho})\,,$ (36)
@@
where $\rho_{\textbf{k}}$ is the Fourier transform of the $k$ mode energy density. In pure kination with initial conditions $\bar{\phi}(t_{0})=0$ and $\dot{\bar{\phi}}(t_{0})=\Pi_{0}$ and the background scalar field evolves as
$\bar{\phi}(t)=\Pi_{0}t_{0}\ln\left(\frac{t}{t_{0}}\right)\,.$ (37)
If we ignore the source term in the perturbation equation of motion, Eq. 34, the background equation can be solved analytically. In deep sub-horizon limit $k\gg H_{0}$ we can obtain an analytical expression with respect to the Bessel functions of zeroth kind:
$\delta\phi_{\textbf{k}}=At^{-\frac{3c-1}{2}}J_{\alpha}\left(\frac{ka^{-1}}{1-c}t\right)+Bt^{-\frac{3c-1}{2}}J_{-\alpha}\left(\frac{ka^{-1}}{1-c}t\right)$ (38)
with
$\alpha=\frac{3c-1}{2c-2}\equiv 0.$ (39)
Then from the asymptotic behaviour of zeroth Bessel functions we obtain
$\delta\phi_{\textbf{k}}\sim\frac{\cos\left(\frac{ka^{-1}}{1-c}\right)}{a}\sim a^{-1}$ (40)
@@
We then conclude that the gradient density of the perturbations evolves as
$\rho_{\text{grad}}\sim\frac{\delta\phi_{\textbf{k}}^{2}}{a^{2}}\sim a^{-4}\,,$ (41)
while the kinetic energy of the perturbations
$\rho_{\text{kin}}\sim\dot{\delta\phi_{\textbf{k}}}^{2}\sim a^{-4}\,,$ (42)
and therefore $\rho_{\delta\phi}\sim a^{-4}$. In the super-horizon regime $k\ll H$ and the gradient term can be ignored.
$\rho_{\delta\phi}\sim\rho_{\text{grad}}\sim\frac{\delta\phi_{\textbf{k}\,0}^{2}}{a^{2}}\sim a^{-2}\,$ (43)
@@
in the limit where the super-horizon modes are frozen. The conclusion is that deep sub-horizon modes grow with respect to the kination background as $a^{4}$ while for super-horizon modes as $a^{2}$.
@@ Appendix B Convergence Tests
We test the robustness of our numerical simulation by comparing the evolution with initial wavelength $\lambda_{0}=20H_{0}^{-1}$ and initial non-linearity parameter $b=0.0005$ in three different base resolutions, namely $N_{LR}=128$, $N_{MR}=144$, $N_{HR}=160$. Figure 7, we show the volume averaged Hamiltonian constraint violation
$\mathcal{H}=R^{3}+K^{2}-K_{ij}K^{ij}-16\pi\rho$ (44)
and the differences between the base resolutions, indicating second order convergence prior to the black hole formation according to Eq. (45).
[…]
@@
Figure 7: Hamiltonian convergence tests for volume-average Hamiltonian constraint violation with initial perturbation characteristic wavelength $\lambda_{0}=20H_{0}^{-1}$, and initial non-linearity parameter $b=0.0005$, Upper Panel: The volume-averaged Hamiltonian constraint with grid resolutions $N_{LR}=128$, $N_{MR}=144$, $N_{HR}=160$, Lower Panel: The difference in volume-averaged Hamiltonian constraint between high-middle and middle-low resolutions, and the second order convergence line from abs(HR-MR) to abs(MR-LR) according to Eq. (45).
[…]

### v2

@@ 4 Implications for Cosmology
$N_{\rm kin}\gtrsim 16.46+\log\left(\frac{10^{14}\mathrm{GeV}}{H_{\rm end}}\right)\,$, (4.1)
@@
$\Delta^{2}_{\mathcal{R}}(k)=\Delta^{2}_{\mathcal{R},\mathrm{CMB}}(k)+\Delta^{2}_{\mathcal{R},\mathrm{peak}}\delta(k-k_{\rm PBH})\,.$ (4.2)
@@
$\beta\sim\frac{\Delta_{\mathcal{R},\mathrm{peak}}}{\sqrt{2\pi}\delta_{c}}\exp(-\frac{\delta_{c}^{2}}{2\Delta^{2}_{\mathcal{R},\mathrm{peak}}})\,.$ (4.3)
@@
$\beta\gtrsim\left(\frac{\Gamma_{\rm ev}}{H_{\rm end}}\right)^{1/2}e^{-\frac{3}{2}N_{\rm kin}+3(N_{\rm form}-N_{\rm end})}\,$, (4.4)
@@
In Figure 7, we translate the results obtained in Section 3 for $\delta_{c}$ into lower bounds on $\Delta^{2}_{\mathcal{R},\mathrm{peak}}$ that satisfy the condition of Eq. (4.4) for an arbitrary value of $H_{\rm end}=10^{13}\mathrm{GeV}$. Nevertheless, due to the logarithmic dependency of $\Delta^{2}_{\mathcal{R},\mathrm{peak}}$ on $\beta$, we note that the results depend only very mildly on the actual value of $H_{\rm end}$. From the figure, it is clear that the very small values of $\delta_{c}$ we obtained for large perturbations (which form at larger $N_{\rm form}$) reduce the required amplitude of the power spectrum by more than an order of magnitude compared to previous claims in the literature that use the analytical results of [31].
@@
This result is particularly important, as it is generally very challenging–from a model-building perspective–to enhance the power spectrum of comoving scalar curvature perturbations all the way up to $\Delta^{2}_{\mathcal{R},\mathrm{peak}}\sim\mathcal{O}(10^{-2}\sim 10^{-3})$ without requiring significant fine-tuning of the inflaton potential parameters [17]. Moreover, such enhancements often risk conflicting with the CMB predictions, especially regarding the spectral index $n_{s}$ and the tensor-to-scalar ratio $r$. However, according to our results, the required power spectrum amplitude during a kination era for PBHs to play a significant cosmological role only needs to reach $\sim\mathcal{O}(10^{-5})$. Achieving this level of enhancement within an inflationary scenario featuring a transient ultra-slow-roll phase does not suffer from the same fine-tuning issues, and is therefore more plausible from an ultraviolet complete theoretical framework.
5 Summary and outlook
@@
We have studied the evolution of scalar inhomogeneities during a kination epoch beyond the linear regime using numerical relativity. Starting with small perturbations, our results in the sub-horizon limit agree with perturbative analysis: linear perturbations grow relative to the background and eventually dominate the universe, but do not collapse into black holes. We then examined initially large perturbations and found that they rapidly begin to behave like radiation unless the initial overdensity is large $\delta_{0}\gg 1$, in which case gravitational forces are strong enough to induce immediate collapse into a black hole. However, such large sub-horizon perturbations cannot arise from the growth of initially small fluctuations after horizon re-entry: their domination over the background would halt further growth, preventing collapse. We therefore conclude that only mechanisms capable of sourcing large sub-horizon inhomogeneities at early times could lead to gravitational collapse.
@@
On the other hand, super-horizon inhomogeneities manifest a much more interesting phenomenology with regard to strong gravity effects. The very long oscillation timescales of these modes imply that they behave like curvature perturbations prior to horizon re-entry. We observed that even in the limit where $\delta_{0}\ll 1$, such perturbations grow significantly relative to the background within just a few e-folds of kination. After horizon re-entry, their fate depends on their amplitude, they may either evolve into a radiation-like background–similar to typical for sub-horizon modes–or collapse into black holes. Our key finding is that the critical overdensity extrapolated from our numerical evolution, is much smaller compared to the perturbative predictions, although it converges to the perturbative predictions when the initial wavelength approaches $H_{0}^{-1}$. For increasingly larger wavelengths, we find that the extrapolated critical overdensity decreases, reaching around $\delta_{c,\text{numerical}}\approx 0.02$ at $\lambda_{0}\approx 2000H_{0}^{-1}$, as shown in Figure 4.
@@
This significant reduction of the extrapolated critical overdensity for the collapse of super-horizon modes into PBHs as compared to previously existing results in the literature is crucial. Indeed, the capacity of Gaussian perturbations—characterized by a primordial power spectrum $\Delta_{\mathcal{R}}^{2}$–to produce PBHs from their horizon re-entry is exponentially sensitive to this critical value, $\beta\simeq\exp(-\delta_{c}^{2}/\Delta_{\mathcal{R}}^{2})$. Reducing $\delta_{c}$ by over an order of magnitude thus considerably eases PBH formation. We estimated the typical value that would be required for $\Delta_{\mathcal{R}}^{2}$ for PBH formation during kination to lead to PBH domination and evaporation before BBN, and found that values of $\mathcal{O}(10^{-5})$ were sufficient–compared to around $\mathcal{O}(10^{-3})$ in the standard kination scenario.
@@
It is also worth noting that in our numerical simulations, the requirement that the initial spatial hyperslice be conformally flat, combined with the imposition of periodic boundary conditions, enforces a high degree of symmetry in our physical setup. Intuitively, one might expect that in the case of black hole formation, the collapse of a single perturbation mode in three spatial directions might remove some energy from the system to create the spherical configuration that eventually collapses. Therefore, we cannot exclude the scenario in which the numerical value of $\delta_{c}$ in the case of direct spherical collapse might be even lower than what we have inferred in section 3. It is important that future simulations explore such spherically symmetric configurations so that the expression for $\delta_{c}$ can be verified more accurately; in this work, we did not pursue this direction due to the difficulty in finding initial conditions that both respect spherical symmetry and satisfy the integrability conditions imposed by periodic boundary conditions. On the opposite, additional degrees of non-sphericity might as well render the collapse more difficult to achieve [escrivà2024simulationsellipsoidalprimordialblack], which also requires a specific study in the context of kination.
@@
Finally, it is necessary to comment on the choice of gauge. In our numerical simulations, we employ a modified version of the moving puncture gauge, which includes a contribution from the volume-averaged extrinsic curvature scalar. As explained in Section 3, this particular gauge choice leads to the lapse function eventually dropping to very low values, effectively freezing the simulations for very large initial scalar wavelengths. A more suitable gauge could enable the evolution of the system for at least seven e-folds of kination, potentially allowing for a more accurate determination of $\delta_{c}$ in the limit of very large initial wavelengths. We leave such explorations to future work.
6 Acknowledgements
We acknowledge useful conversations with Martin Mosny, Joseph Conlon, Benjamin Stefanek, and Katy Clough. CC and PG thank Santiago Agüí Salcedo, Gonzalo Villa, Lucas Vicente Garcia-Consuegra, Josu Calvo Aurrekoetxea, Christian Byrnes, Robyn Munoz, Daniel G. Figueroa and Liina Chung-Jukko for their useful input. We would also like to thank the GRChombo team (http://www.grchombo.org/). This work is supported by a Research Project Grant RPG-2021-423 from Leverhulme Trust.
@@ Appendix A Linear Perturbations During Kination
We briefly review the relevant solutions for scalar perturbations during kination on a fixed FLRW background. For a more detailed analysis, including the effects of metric perturbations, we refer the reader to [5, eröncel2025universalbounddurationkination, 36]. The scalar field in the presence of small perturbations can be written as
$\phi(\eta,\vec{x})=\bar{\phi}(\eta)+\delta\phi(\eta,\vec{x})=\bar{\phi}(\eta)+\frac{f(\eta,\vec{x})}{a}\,.$ (A.1)
The perturbation can be decomposed into Fourier modes as
$\delta\phi(\eta,\vec{x})=\int\frac{d^{3}k}{(2\pi)^{3}}\delta\phi_{k}(\eta)e^{i\vec{k}\cdot\vec{x}}\,.$ (A.2)
@@
The equations of motion for the background scalar field $\bar{\phi}$ and its perturbations in Fourier space, in the limit where the scalar potential vanishes ($V\to 0$), are given by
${\bar{\phi}}^{\prime\prime}+2\mathcal{H}{\bar{\phi}}^{\prime}=0\,$, (A.3)
$f_{k}^{\prime\prime}+\left(k^{2}-\frac{a^{\prime\prime}}{a}\right)f_{k}=0\,$, (A.4)
where the perturbation equation is the Mukhanov–Sasaki equation. For a kination background, this equation becomes
$f_{k}^{\prime\prime}+\left(k^{2}+\frac{1}{4\eta^{2}}\right)f_{k}=0\,.$ (A.5)
During kination, the scale factor evolves as
$a=a_{0}\left(\frac{\eta}{\eta_{0}}\right)^{1/2}\,$, (A.6)
and the conformal Hubble parameter is
$\mathcal{H}=\frac{1}{2\eta}\,$, (A.7)
while the background density evolves as
$\bar{\rho}\sim a^{-6}\,.$ (A.8)
@@
In pure kination, assuming initial conditions $\bar{\phi}(\eta_{0})=0$ and $\bar{\phi}^{\prime}(\eta_{0})=\Pi_{0}$, the background scalar field evolves as
$\bar{\phi}(\eta)\sim\ln\left(\frac{\eta}{\eta_{0}}\right)\,.$ (A.9)
The general solution of the Mukhanov–Sasaki equation can be expressed in terms of Bessel functions $J_{0}$ and $Y_{0}$,
$f_{k}(\eta)=A\sqrt{\eta}\,J_{0}(k\eta)+B\sqrt{\eta}\,Y_{0}(k\eta)\,.$ (A.10)
We now investigate the asymptotic limits of this solution.
In the deep sub-horizon limit, $k\eta\gg 1$, the asymptotic behavior of the Bessel functions leads to oscillatory solutions,
$\delta\phi_{k}=\frac{f_{k}}{a}\sim\frac{\cos(k\eta)}{a}\sim a^{-1}\,.$ (A.11)
@@
We therefore conclude that the gradient energy density of the perturbations evolves as
$\langle\rho_{\nabla}\rangle\sim\left\langle\frac{\delta\phi_{k}^{2}}{a^{2}}\right\rangle\sim a^{-4}\,$, (A.12)
while the kinetic energy density of the perturbations evolves as
$\langle\rho_{\delta K}\rangle$ $=\left\langle\left(\frac{f_{k}^{\prime}}{a^{3}}-\mathcal{H}\frac{f_{k}}{a^{2}}\right)\bar{\phi}^{\prime}\right\rangle$
$\quad+\left\langle\left(\frac{f_{k}^{\prime 2}}{2a^{4}}+\mathcal{H}^{2}\frac{f_{k}^{2}}{2a^{2}}-\mathcal{H}\frac{f_{k}^{\prime}f_{k}}{a^{3}}\right)\right\rangle\sim a^{-4}\,.$ (A.13)
Therefore, the total perturbation energy density evolves as
$\rho_{\delta\phi}=\langle\rho_{\nabla}+\rho_{\delta K}\rangle\sim a^{-4}\,.$ (A.14)
The density contrast in the sub-horizon regime is therefore
$\delta_{\text{sub}}=\frac{\rho_{\delta\phi}}{\bar{\rho}}\sim a^{2}\,.$ (A.15)
In the super-horizon regime, $k\eta\ll 1$, the solution becomes
$f_{k}=c_{1}\eta^{1/2}+c_{2}\eta^{1/2}\ln\eta\quad\Rightarrow\quad f_{k}=Aa+Ba\ln a\,.$ (A.16)
Consequently,
$\delta\phi_{k}=A+B\ln a\,.$ (A.17)
Note that, unlike the de Sitter case where $a^{\prime\prime}/a=2/\eta^{2}$ and the perturbations freeze completely outside the horizon, the scalar perturbations during kination evolve logarithmically. However, if the initial velocity satisfies $\delta\phi^{\prime}(\eta_{0})=0$, then $B=0$, eliminating the logarithmic contribution, and the mode remains frozen while outside the horizon.
In this case,
$\rho_{\delta\phi}\sim\langle\rho_{\nabla}\rangle\sim\frac{\delta\phi_{k,0}^{2}}{a^{2}}\sim a^{-2}\,.$ (A.18)
@@
The conclusion is that deep sub-horizon modes grow relative to the kination background as
$\delta_{\text{sub}}\sim a^{2}\,$, (A.19)
while super-horizon modes grow as
$\delta_{\text{super}}\sim a^{4}\,.$ (A.20)
If super-horizon perturbations in the kinetic sector are also included, the logarithmic term contributes nontrivially. However, we find that
$\langle\rho_{\delta K}\rangle\sim a^{-6}\,$, (A.21)
which means that the kinetic perturbations decay at the same rate as the background. Therefore, the dominant contribution comes from the gradient energy density, which decays in the same way as scalar curvature perturbations.
@@ Appendix B Convergence Tests
We test the robustness of our numerical simulation by comparing the evolution with initial wavelength $\lambda_{0}=20H_{0}^{-1}$ and initial density contrast $\delta_{0}=0.0005$ in three different base resolutions, namely $N_{LR}=128$, $N_{MR}=144$, $N_{HR}=160$. Figure 8, we show the volume averaged Hamiltonian constraint violation
$\mathcal{H}=R^{3}+K^{2}-K_{ij}K^{ij}-16\pi\rho$ (B.1)
and the differences between the base resolutions, indicating second order convergence prior to the black hole formation according to Eq. (B.2).
$\frac{\|{\langle\mathcal{H}_{MR}\rangle-\langle\mathcal{H}_{LR}\rangle}\|}{\|{\langle\mathcal{H}_{HR}\rangle-\langle\mathcal{H}_{MR}\rangle}\|}\approx\frac{1/N_{MR}^{p}-1/N_{LR}^{p}}{1/N_{HR}^{p}-1/N_{MR}^{p}}$ (B.2)
@@
Figure 8: Hamiltonian convergence tests for volume-average Hamiltonian constraint violation with initial perturbation characteristic wavelength $\lambda_{0}=20H_{0}^{-1}$, and initial non-linearity parameter $\delta_{0}=5\times 10^{-4}$, Upper Panel: The volume-averaged Hamiltonian constraint with grid resolutions $N_{LR}=128$, $N_{MR}=144$, $N_{HR}=160$, Lower Panel: The difference in volume-averaged Hamiltonian constraint between high-middle and middle-low resolutions, and the second order convergence line from abs(HR-MR) to abs(MR-LR) according to Eq. (B.2), showing good level of second order convergence.
Figure 9: Convergence tests for black hole mass formed with initial perturbation characteristic wavelength $\lambda_{0}=20H_{0}^{-1}$, and initial density contrast $\delta_{0}=5\times 10^{-4}$, Upper Panel: The measured black hole mass with apparent horizon finder with grid resolutions $N_{LR}=128$, $N_{MR}=144$, $N_{HR}=160$, Lower Panel: Errors in black hole mass between high-middle and middle-low resolutions showing convergence to $5\%$.